How do you solve this problem: The square route of k+9 minus the square route of k equals 3.
Answer by stanbon(57245)

The square route of k+9 minus the square route of k equals 3.
--------
It's not square route; it's square root.
----------
sqrt(k+9) - sqrt(k) = 3
Isolate one of the sqrt terms, as follows:
sqrt(k+9)= sqrt(k) + 3
Square both sides to get:
k+9 = k+6sqrtk+9
6sqrt(k)=0
sqrtk=0
k=0
----------
Check that answer in the original equation:
sqrt(0+9) -sqrt(0)= 3
sqrt(0) - 0 = 3
3=3
It checks out alright.
Cheers,
Stan H.